OpenStudy (anonymous):

Algebra 2 Work Check

OpenStudy (anonymous):

\[f(x)\frac{ x+7 }{ 5 } \] solve for \[f^-1(3) \] I have a answer of \[f(3)=\frac{ 3+7 }{ 5 }\]

OpenStudy (anonymous):

@agent0smith @SolomonZelman

OpenStudy (agent0smith):

You need to find the inverse function first, THEN plug in 3. You plugged 3 into the original function.

OpenStudy (anonymous):

so it would be x=y+7/5 then what